OpenStudy (anonymous):

f(x)=[lnx]4 f'x and f(e^3)'

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

ln(x) to the 4th power?

OpenStudy (anonymous):

are you saying find f'(x)? if so then f(x)=4[ln(x)] f'(x)= 4(1/x) f'(x)= 4/x and to find f(e^3)' just plug that into =4/(e^3)

OpenStudy (anonymous):

yes 4th power

jimthompson5910 (jim_thompson5910):

you have to use the chain rule to get f(x) = [ ln(x) ]^4 f'(x) = 4[ ln(x) ]^3 * d/dx[ ln(x) ] f'(x) = 4[ ln(x) ]^3 * (1/x) then you plug in x = e^3 and evaluate

OpenStudy (anonymous):

Wow, I see its so clear! People gave me answers but none of them were corrrect you did it! cool!

OpenStudy (anonymous):

f(x)=5cos(7ln(x)) , find f(x

OpenStudy (anonymous):

this is the next question but i will post it to new one so you could have medal